OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth 1)11 2)7.1 3)7 4)-7.1

OpenStudy (amistre64):

so many homework questions, and not enough time to do them ;)

OpenStudy (anonymous):

im am doing them , i just need more help on these.

OpenStudy (anonymous):

\[\sqrt{(-2-(-7)^2+(-3-(-8))^2}\] \[\sqrt{5^2+5^2}\] \[\sqrt{50}\]

OpenStudy (amistre64):

i like to stack the points; makes subtracting them easier

OpenStudy (anonymous):

i get 7.07 rounds to 7.1

OpenStudy (amistre64):

( -2, -3) -(-7, -8) ( -2, -3) ( 7, 8) --------- (5,5) sqrt(5^2 + 5^2) = sqrt(50) = sqrt(2*25) = 5sqrt(2)

OpenStudy (anonymous):

another embarrassment frankly since it is really \[5\sqrt{2}\] as you have the diagonal of a square with side 5!